OpenStudy (anonymous):

the gas tank of a truck is a cylinder 5ft long with a diameter of 1.5ft. at the gas station, a pump pours gas at the rate of 2 ft^3 per minute. how many minutes does it take to fill the empty tank with that pump? use the value 3.14 for pi

OpenStudy (anonymous):

do you know how to get the area of a circle? and hence the volume of a cylinder?

OpenStudy (anonymous):

no

OpenStudy (anonymous):



OpenStudy (anonymous):

can you help me doing this problem step by step

OpenStudy (anonymous):

so that is what you have been given. You first need to find the volume of a cylinder. \[v=h*\pi*r^2\] \[r=\frac{1}{2}*D\] so you have \[h=5, D=1.5 \] and thus \[r=0.75\]

OpenStudy (anonymous):

you should now be able to work out v

OpenStudy (anonymous):

would this be the next step 3.14 * 0.75^2 * 3.5 = 6.19

OpenStudy (anonymous):

is h supposed to be 3.5? in your question it says 5

OpenStudy (anonymous):

so it would be 3.14*0.75^2*5=8.83125 then what would be the next step?

OpenStudy (anonymous):

yes that is the volume of the cylinder

OpenStudy (anonymous):

okay is there a next step?

OpenStudy (anonymous):

So the next step is to find how long it takes to fill. You have been given that that the pump pours \(2\)ft\(^3\) per minute. This is saying that every minute the volume of the cylinder has been filled by \(2\)ft\(^3\). there for you need to now calclulate \[\frac{v}{2}\]

OpenStudy (anonymous):

would it be this 4.415625

OpenStudy (anonymous):

yes. a good trick to do with your calculator is there is a button that looks like.....

OpenStudy (anonymous):